# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

### v2

@@ B.3 Proof of Proposition 1
Condition (a) readily fixes the signs of the columns of $B_{1},...,B_{M}$, so it suffices the check that under Conditions (c) and (b), changing the ordering of the columns of any of $B_{2},...,B_{M}$ would change the joint density of the model. Suppose the observed time series is $y_{t-p+1},...,y_{0},y_{1},...,y_{T}$ and denote $Y_{T}=(y_{T},...,y_{0})$. Conditional on the initial values $y_{t-p+1},...,y_{0}$, the joint density of the model is
$f_{Y}(Y_{T})=\prod_{t=1}^{T}\left(|\det(B_{t})|^{-1}\prod_{i=1}^{d}d_{e_{i}}(I_{d,i}^{\prime}B_{y,t}^{-1}(y_{t}-\mu_{t});0,1,\nu_{i})\right),$ (B.5)
@@
where $I_{d,i}$ is the $i$th column of the $d$-dimensional identity matrix and $d_{e_{i}}(\cdot;0,1,\nu_{i})$ is the density function the $i$th shock. At each $t$, $B_{y,t}=B_{m}$ for the regime $m$ with $\alpha_{m,t}=1$, implying the that the density (B.5) can be expressed as
$f_{Y}(Y_{T};B_{1},...,B_{M})$ $=\prod_{m=1}^{M}\prod_{t\in\mathcal{T}_{m}}f_{y}(y_{t}-\mu_{t};B_{m}),\ \ \text{and}$ (B.6)
$f_{y}(y_{t}-\mu_{t};B_{m})$ $=|\det(B_{m})|^{-1}\prod_{i=1}^{d}d_{e_{i}}(I_{d,i}^{\prime}B_{m}^{-1}(y_{t}-\mu_{t});0,1,\nu_{i}),$ (B.7)
@@
where $\mathcal{T}_{m}$, $m=1,...,M$, is the (nonempty) set of time indices that satisfy $\alpha_{m,t}=1$ for the regime $m$, and the dependence of the density functions $f_{Y}(\cdot;B_{1},...,B_{M})$ and $f_{y}(\cdot;B_{m})$ on the impact matrices $B_{1},...,B_{M}$ is now made explicit in the notation (but its dependence on the other parameters is omitted).
@@
Denote by $\boldsymbol{\tau}_{m}=(\tau_{m1},...,\tau_{md})$ the permutation $\tau_{m1},...,\tau_{md}$ of the columns of $B_{m}$, and denote by $B_{m}(\boldsymbol{\tau}_{m})$ the impact matrix $B_{m}$ corresponding to the permutation $\boldsymbol{\tau}_{m}$ of its columns. Then, consider the permutations $\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{2},...,\boldsymbol{\tau}_{M}$ and $\boldsymbol{\tau}_{1},\tilde{\boldsymbol{\tau}}_{2},...,\tilde{\boldsymbol{\tau}}_{M}$. By the invertibility of $B_{m}$, none of the entries in $B_{m}^{-1}(y_{t}-\mu_{t})$ are equal to each other almost everywhere in $Y_{T}\in\mathbb{R}^{T}$. Since reordering the columns of $B_{m}$ reorders the rows of $B_{m}^{-1}$, it then follows that if $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$, some of the entries in $B_{m}^{-1}(\boldsymbol{\tau}_{m})(y_{t}-\mu_{t})$ and $B_{m}^{-1}(\tilde{\boldsymbol{\tau}}_{m})(y_{t}-\mu_{t})$ are in a different order. Moreover, as $|\det(B_{m})|$ is invariant to the ordering of the columns of $B_{m}$, it follows from this and Condition (c) that, almost everywhere in $Y_{T}\in\mathbb{R}^{T}$, $f_{y}(y_{t}-\mu_{t};B_{m}(\boldsymbol{\tau}_{m}))\neq f_{y}(y_{t}-\mu_{t};B_{m}(\tilde{\boldsymbol{\tau}}_{m}))$ whenever $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$.
Therefore, as the shocks and thereby $y_{t}$ follows a continuous distribution, we have
@@
almost everywhere in $Y_{T}\in\mathbb{R}^{T}$ whenever $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$ for any $m=2,...,M$. Since the this result holds almost everywhere in $Y_{T}\in\mathbb{R}^{T}$, and $B_{y,t}$ is here invertible for all invertible $B_{1},...,B_{M}$, the result holds almost everywhere in $Y_{T}\in\mathbb{R}^{T}$.
Finally, by Lemma 2, $B_{y,t}$ is identified up to ordering and signs of its columns at each $t$. Thus, each impact matrix $B_{1},...,B_{M}$ of the regimes is identified up to ordering signs of its columns. Under the conditions of Proposition 1, the signs of the columns of $B_{1},...,B_{M}$ and the ordering of the columns $B_{1}$ are directly fixed. Moreover, by the above discussion, different permutations of the columns of any of $B_{2},...,B_{M}$ lead to different values of the joint density function of the model almost everywhere in $Y_{T}\in\mathbb{R}^{T}$, i.e., to models that are not observationally equivalent. Thus, only models corresponding to the same permutations of the columns of $B_{2},...,B_{M}$ are observationally equivalent, implying that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $Y_{T}\in\mathbb{R}^{T}$.∎
@@ B.4 Proof of Proposition 2
To proof consists of two parts. First, we show that almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, if the ordering and signs of the columns of $B_{1}$ is fixed, changing ordering or signs of the columns of any of $B_{2},...,B_{M}$ leads to an impact impact matrix $B_{y,t}$ that cannot, at all $t$, be obtained by reordering or changing the signs of the columns of the original impact matrix. In other words, we show that fixing the ordering signs of the columns of $B_{1}$ effectively fixes the ordering and signs of the columns of $B_{2},...,B_{M}$ and thereby of $B_{y,t}$. Then, we make use of this result together with the result of Lemma 2 (that $B_{y,t}$ is identified up ordering and signs of its columns at each $t$) to show that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$.
Denote by $\boldsymbol{\tau}_{m}=(\tau_{m1},...,\tau_{md})$ the permutation $\tau_{m1},...,\tau_{md}$ and $\boldsymbol{\iota}_{m}=(\iota_{m1},...,\iota_{md})$ the sign changes $\iota_{m1},...,\iota_{md}$ of the columns of $B_{m}$, respectively. Moreover, denote by $B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})$ the invertible impact matrix $B_{m}$ corresponding to the permutation $\boldsymbol{\tau}_{m}$ and sign changes $\boldsymbol{\iota}_{m}$ of its columns. Then, consider the permutations $\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{2},...,\boldsymbol{\tau}_{M}$ and $\boldsymbol{\tau}_{1},\tilde{\boldsymbol{\tau}}_{2},...,\tilde{\boldsymbol{\tau}}_{M}$ as well the sign changes $\boldsymbol{\iota}_{1},\boldsymbol{\iota}_{2},...,\boldsymbol{\iota}_{M}$ and $\boldsymbol{\iota}_{1},\tilde{\boldsymbol{\iota}}_{2},...,\tilde{\boldsymbol{\iota}}_{M}$. We start by showing that $B_{y,t}(\boldsymbol{\tau},\boldsymbol{\iota})\equiv\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})$ cannot be obtained by reordering or changing the signs of the columns of $B_{y,t}(\tilde{\boldsymbol{\tau}},\tilde{\boldsymbol{\iota}})\equiv\alpha_{m,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m})$ almost everywhere in $[B_{1},...,B_{M}]\in\mathbb{R}^{d\times Md}$ whenever $\boldsymbol{\tau}_{m}\neq\tilde{\boldsymbol{\tau}}_{m}$ or $\boldsymbol{\iota}_{m}\neq\tilde{\boldsymbol{\iota}}_{m}$ for any $m=2,...,M$.
@@
By the above discussion, for any single transformation $D_{t}P_{t}$ of $B_{y,t}$, the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (B.11) for any time period when all the entries of $(\alpha_{1,t},...,\alpha_{M,t})$ are nonzero has a Lebesgue measure zero in $[B_{1}:\cdots B_{M}]\in\mathbb{R}^{d\times Md}$. Since there exists a finite number of such transformations and a countable union of measure zero sets has a measure zero, it follows that the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (B.12) for all $t$ has a Lebesgue measure zero in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$.
@@
Under Condition (C), the vector $(\alpha_{1,t},...,\alpha_{M,t})$ and all its subvectors $(\alpha_{n_{1},t},...,\alpha_{n_{M^{\circ}},t})$, $n_{l}\in\mathbb{M^{\circ}}$, $\mathbb{M^{\circ}}\subset\{1,...,M\}$, where $M^{\circ}\geq 2$ denotes the number of elements in $\mathbb{M^{\circ}}$, take at least $M+1$ linearly independent values (plus one accounting for transition weights that equal to zero). Therefore, Equation (B.14) defined together for all $t$ defines a system of linear equations involving $Md^{2}$ variables and at least $Md^{2}$ linearly independent equations. Moreover, as $(\alpha_{1,t},...,\alpha_{M,t})$ and all its subvectors take at least $M+1$ linearly independent values, each variable $\tilde{b}_{ij}^{(m)}$ appears in at least $M$ linearly independent the equations. Thus, there exists at most one solution to this system of linear equations. Since one solution is $\tilde{B}_{m}=B_{m}$ for all $m=1,...,M$, this is the unique solution, concluding that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywehere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$.∎

### v3

@@ B.3 Proof of Proposition 1
First, we establish the following lemma and its corollary, the latter of which is used in the proof:
Lemma B.2.
Suppose $(\nu_{i},\lambda_{i})\neq(\nu_{j},\lambda_{j})$ (Condition (b) of Proposition 1). Then, $st(x_{1};\nu_{i},\lambda_{i})\neq st(x_{2};\nu_{j},\lambda_{j})$ almost everywhere in $(x_{1},x_{2})\in\mathbb{R}^{2}$, where the density function $st(\cdot;\nu_{i},\lambda_{i})$ is given in Equation (4.1). Moreover, if $\lambda_{i}\neq 0$, then the equation $st(x;\nu_{i},\lambda_{i})=st(-x;\nu_{i},\lambda_{i})$ has at most six solutions in $x\in\mathbb{R}$, and thus, $st(x;\nu_{i},\lambda_{i})\neq st(-x;\nu_{i},\lambda_{i})$ almost everywhere in $x\in\mathbb{R}$.
Proof.
Denote $f(x_{1},x_{2})\equiv st(x_{1};\nu_{i},\lambda_{i})-st(x_{2};\nu_{j},\lambda_{j})$. It is straightforward to show (by induction) that the function $st(x_{i};\nu_{i},\lambda_{i})$, and also thereby $f(x_{1},x_{2})$, is infinitely continuously differentiable. It then follows from Implicit function theorem that in the points $x^{0}=(x_{1,0},x_{2,0})\in\mathbb{R}^{2}$ where the gradient of $f(x_{1},x_{2})$ is nonzero, the set $\mathcal{X}\equiv\{(x_{1},x_{2})\in\mathbb{R}^{2}|f(x_{1},x_{2})=0\}$ is an one-dimensional manifold locally in a neighbourhood of $x^{0}$. Because this argument applies at all points $x^{0}$ where the gradient of $f(x_{1},x_{2})$ is nonzero, it follows that the set $\mathcal{X}$ is a one-dimensional manifold in regions where this condition holds. Therefore, Lebesque measure of the set $\mathcal{X}$ is zero in $\mathbb{R}^{2}$ in the region where the gradient of $f(x_{1},x_{2})$ is nonzero. On the other hand, by Sard’s theorem (Sard, 1994), the set of points where the gradient of $f(x_{1},x_{2})$ is zero has a Lebesgue measure zero in $\mathbb{R}^{2}$. Since the union of two sets with a Lebesgue measure zero has a Lebesgue measure zero, these results can be combined to conclude that the $\mathcal{X}$ has a Lebesgue measure zero in $\mathbb{R}^{2}$. That is, $st(x_{1};\nu_{i},\lambda_{i})=st(x_{2};\nu_{j},\lambda_{j})$ only on sets of zero Lebesgue measure, concluding that $st(x_{1};\nu_{i},\lambda_{i})\neq st(x_{2};\nu_{j},\lambda_{j})$ almost everywhere in $(x_{1},x_{2})\in\mathbb{R}^{2}$.
To prove the moreover part, suppose $\lambda_{i}\neq 0$ and $st(x;\nu_{i},\lambda_{i})=st(-x;\nu_{i},\lambda_{i})$, which simplifies to
$\left(\frac{b_{i}x+a_{i}}{1-\mathbbm{1}\{x<-a_{i}/b_{i}\}\lambda_{i}+\mathbbm{1}\{x\geq-a_{i}/b_{i}\}\lambda_{i}}\right)^{2}=\left(\frac{-b_{i}x+a_{i}}{1-\mathbbm{1}\{-x<-a_{i}/b_{i}\}\lambda_{i}+\mathbbm{1}\{-x\geq-a_{i}/b_{i}\}\lambda_{i}}\right)^{2},$ (B.12)
where $a_{i}\neq 0$ and $b_{i}\neq 1$. If $x<-a_{i}/b_{i}$ and $-x<-a_{i}/b_{i}$ or $x\geq-a_{i}/b_{i}$ and $-x\geq-a_{i}/b_{i}$, a direct calculation shows that the only solution to Equation (B.12) is $x=0$. If $x<-a_{i}/b_{i}$ and $-x\geq-a_{i}/b_{i}$ or $x\geq-a_{i}/b_{i}$ and $-x<-a_{i}/b_{i}$, Equation (B.12) is a second degree polynomial equation in $x$ and has thereby at most two real solutions. Combining all the four possible cases, in can be concluded that $st(x;\nu_{i},\lambda_{i})=st(-x;\nu_{i},\lambda_{i})$ has at most six solutions in $x\in\mathbb{R}$. ∎
Corollary B.1.
Suppose $D$ is a diagonal matrix with ±$\pm 1$ diagonal entries and $P$ is a permutation matrix, and denote by $I_{d,i}$ the $i$th column of the $d$-dimensional identity matrix. Moreover, suppose the entries $x$ are distinct as is and after changing any of their signs. If $(\nu_{i},\lambda_{i})\neq(\nu_{j},\lambda_{j})$ for all $i\neq j=1,...,d$ and $\lambda_{i}\neq 0$ for all $i=1,...,d$, and in addition, $D\neq I_{d}$ or $P\neq I_{d}$, then
$\prod_{i=1}^{d}st(I_{d,i}^{\prime}x;\nu_{i},\lambda_{i})\neq\prod_{i=1}^{d}st(I_{d,i}^{\prime}DPx;\nu_{i},\lambda_{i})$ (B.13)
almost everywhere in $x\in\mathbb{R}^{d}$.
Proof.
Suppose $\prod_{i=1}^{d}st(I_{d,i}^{\prime}x;\nu_{i},\lambda_{i})=\prod_{i=1}^{d}st(I_{d,i}^{\prime}DPx;\nu_{i},\lambda_{i})$. Rearranging and taking logarithm yields
$\sum_{i=1}^{d}(\log st(I_{d,i}^{\prime}x;\nu_{i},\lambda_{i})-\log st(I_{d,i}^{\prime}DPx;\nu_{i},\lambda_{i}))=0.$ (B.14)
Since $DPx$ reorders or swaps the signs of some of the elements of $x$ (but does not leave all the elements unchanged), by Lemma B.2 (and strict monotonicity of logarithm),
$f_{i}(x)\equiv\log st(I_{d,i}^{\prime}x;\nu_{i},\lambda_{i})-\log st(I_{d,i}^{\prime}DPx;\nu_{i},\lambda_{i})\neq 0$ (B.15)
for at least one $i\in\{1,...,d\}$ over sets of positive Lebesgue measure. Therefore, Equation (B.14) can hold almost everywhere only if the nonzero terms in the sum cancel each other out.
Denote $\mathcal{I}(x)\equiv\{i\in\{1,...,d\}|f_{i}(x)\neq 0\}$ and $g(x^{*})\equiv\sum_{i\in\mathcal{I}}f_{i}(x)$, where $x^{*}$ is the $(d^{*}\times 1)$-dimensional ($d^{*}\leq d$) subvector of $x$ containing the components of $x$ for which $I_{d,i}^{\prime}x\neq\pm I_{d,i}^{\prime}DPx$, $i=1,...,d$ (i.e., $x^{*}$ captures all variables contributing to any nonzero $f_{i}(x)$). It is straightforward to show (by induction) that the function $\log st(x_{i};\nu_{i},\lambda_{i})$, $x_{i}\in\mathbb{R}$, is infinitely continuously differentiable. It follows that also $g(x^{*})$ is infinitely continuously differentiable, and thus $d^{*}$ times continuously differentiable. Then, arguments based on Implicit function theorem and Sard’s theorem (Sard, 1994) can be used similarly to the proof of Lemma B.2 to conclude that the set of solutions to $g(x^{*})=0$ has a Lebesgue measure zero in $x^{*}\in\mathbb{R}^{d^{*}}$. Hence, as sets of zero Lebesgue measure in $\mathbb{R}^{d^{*}}\subset\mathbb{R}^{d}$ have zero Lebesgue measure in $\mathbb{R}^{d}$, Equation (B.14) can hold only in sets of zero Lebesgue measure in $x\in\mathbb{R}^{d}$, concluding the proof. ∎
@@
Condition (a) of Proposition 1 fixes the ordering and signs of the columns of $B_{1}$, so it suffices to show that under Conditions (b) and (c), changing the ordering or signs of the columns of any of $B_{2},...,B_{M}$ would change the joint density of the model. Suppose the observed time series is $y_{t-p+1},...,y_{0},y_{1},...,y_{T}$ and denote $Y_{T}=(y_{T},...,y_{0})$. Conditional on the initial values $y_{t-p+1},...,y_{0}$, the joint density of the model is
$f_{Y}(Y_{T})=\prod_{t=1}^{T}\left(|\det(B_{y,t})|^{-1}\prod_{i=1}^{d}st(I_{d,i}^{\prime}B_{y,t}^{-1}(y_{t}-\mu_{y,t});\nu_{i},\lambda_{i})\right),$ (B.16)
@@
where $I_{d,i}$ is the $i$th column of the $d$-dimensional identity matrix and $st(\cdot;\nu_{i},\lambda_{i})$ is the density function the $i$th shock given in Equation (4.1). At each $t$, $B_{y,t}=B_{m}$ for the regime $m$ with $r_{m-1}<z_{t}\leq r_{m}$, where the threshold parameters, as well as $B_{y,t}$ up to ordering and signs of its columns, are identified by Lemma 2 regardless of whether $B_{1},...,B_{M}$ are identified. Consequently the density (B.16) can be expressed as
$f_{Y}(Y_{T};B_{1},...,B_{M})$ $=\prod_{m=1}^{M}\prod_{t\in\mathcal{T}_{m}}f_{y}(y_{t}-\mu_{y,t};B_{m}),\ \ \text{and}$ (B.17)
$f_{y}(y_{t}-\mu_{y,t};B_{m})$ $=|\det(B_{m})|^{-1}\prod_{i=1}^{d}st(I_{d,i}^{\prime}B_{m}^{-1}(y_{t}-\mu_{y,t});\nu_{i},\lambda_{i}),$ (B.18)
@@
where $\mathcal{T}_{m}$, $m=1,...,M$, is the (nonempty) set of time indices (identified by Lemma 2) that satisfy $r_{m-1}<z_{t}\leq r_{m}$ for the regime $m$, $\mu_{y,t}$ is the conditional mean of the model (identified by Lemma 2), and the dependence of the density functions $f_{Y}(\cdot;B_{1},...,B_{M})$ and $f_{y}(\cdot;B_{m})$ on the impact matrices $B_{1},...,B_{M}$ is now made explicit in the notation (but their dependence on the other parameters is omitted).
Denote by $\boldsymbol{\tau}_{m}=(\tau_{m1},...,\tau_{md})$ the permutation $\tau_{m1},...,\tau_{md}$ of the columns of $B_{m}$, and denote by $\boldsymbol{\iota}_{m}=(\iota_{m1},...,\iota_{md})$ the sign changes $\iota_{m1},...,\iota_{md}$ of the columns of $B_{m}$. Moreover, denote by $B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})$ the impact matrix $B_{m}$ corresponding to the permutation $\boldsymbol{\tau}_{m}$ and sign changes $\boldsymbol{\iota}_{m}$ of its columns. Then, consider the permutations $\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{2},...,\boldsymbol{\tau}_{M}$ and $\boldsymbol{\tau}_{1},\tilde{\boldsymbol{\tau}}_{2},...,\tilde{\boldsymbol{\tau}}_{M}$ as well as the sign changes $\boldsymbol{\iota}_{1},\boldsymbol{\iota}_{2},...,\boldsymbol{\iota}_{M}$ and $\boldsymbol{\iota}_{1},\tilde{\boldsymbol{\iota}}_{2},...,\tilde{\boldsymbol{\iota}}_{M}$.
@@
By the invertibility of $B_{m}$, almost everywhere in $y_{t}\in\mathbb{R}^{d}$, none of the entries in $B_{m}^{-1}(y_{t}-\mu_{y,t})$ are equal to each other as is nor after changing the signs of some of the entries. Since reordering the columns of $B_{m}$ reorders the rows of $B_{m}^{-1}$, it follows that if $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$, some of the entries in $B_{m}^{-1}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})(y_{t}-\mu_{y,t})$ and $B_{m}^{-1}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m})(y_{t}-\mu_{y,t})$ are in a different order. Likewise, changing the signs of the columns of $B_{m}$ changes the signs of the rows of $B_{m}^{-1}$, it follows that if $\tilde{\boldsymbol{\iota}}_{m}\neq\boldsymbol{\iota}_{m}$, some of the entries in $B_{m}^{-1}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})(y_{t}-\mu_{y,t})$ and $B_{m}^{-1}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m})(y_{t}-\mu_{y,t})$ have different signs (in addition to possibly being in a different order).
By Corollary B.1, different orderings and signs of the elements of $B_{m}^{-1}(y_{t}-\mu_{y,t})$ yield different values of $\prod_{i=1}^{d}st(I_{d,i}^{\prime}B_{m}^{-1}(y_{t}-\mu_{y,t});\nu_{i},\lambda_{i})$ almost everywhere in $y_{t}\in\mathbb{R}^{d}$ under Conditions (b) and (c) of Proposition 1. Moreover, as $|\det(B_{m})|$ is invariant to the ordering and signs of the columns of $B_{m}$, it follows from that, $f_{y}(y_{t}-\mu_{y,t};B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m}))\neq f_{y}(y_{t}-\mu_{y,t};B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m}))$ almost everywhere in $y_{t}\in\mathbb{R}^{d}$ whenever $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$ or $\tilde{\boldsymbol{\iota}}_{m}\neq\boldsymbol{\iota}_{m}$. Therefore, as the shocks and thereby $y_{t}$ follows a continuous distribution, we have
@@
almost everywhere in $Y_{T}\in\mathbb{R}^{dT}$ whenever $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$ or $\tilde{\boldsymbol{\iota}}_{m}\neq\boldsymbol{\iota}_{m}$ for any $m=2,...,M$. Since this result holds almost everywhere in $Y_{T}\in\mathbb{R}^{T}$, and $B_{y,t}$ is here invertible for all invertible $B_{1},...,B_{M}$, the result holds almost everywhere in $Y_{T}\in\mathbb{R}^{dT}$.
By Lemma 2, $B_{y,t}$ is identified up to ordering and signs of its columns at each $t$. Thus, each impact matrix $B_{1},...,B_{M}$ of the regimes is identified up to ordering signs of its columns. Under the conditions of Proposition 1, the ordering and signs of the columns of $B_{1}$ are directly fixed. Moreover, by the above discussion, different permutations or sign changes of the columns of any of $B_{2},...,B_{M}$ lead to different values of the joint density function of the model almost everywhere in $Y_{T}\in\mathbb{R}^{dT}$, i.e., to models that are not observationally equivalent. Thus, only models corresponding to the same permutations and sign changes of the columns of $B_{2},...,B_{M}$ are observationally equivalent, implying that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $Y_{T}\in\mathbb{R}^{dT}$.∎
@@ B.4 Proof of Proposition 2
To proof consists of two parts. First, we show that if the ordering and signs of the columns of $B_{1}$ is fixed, changing the ordering or signs of the columns of any of $B_{2},...,B_{M}$ leads to an impact impact matrix $B_{y,t}$ that cannot, at all $t$, be obtained by reordering or changing the signs of the columns of the original impact matrix. In other words, we show that fixing the ordering signs of the columns of $B_{1}$ effectively fixes the ordering and signs of the columns of $B_{2},...,B_{M}$ and thereby of $B_{y,t}$. Then, we make use of this result together with the result of Lemma 2 (that $B_{y,t}$ is identified up ordering and signs of its columns at each $t$) to show that the matrices $B_{1},...,B_{M}$ are uniquely identified. For models with logistic transition weights (2.4), the result holds almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$, and $z_{t}\in\mathbb{R}^{T}$ (where $z_{t}$ is the switching variable), whereas for models with nonrandom exogenous transition weights, the result holds almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.
By Lemma 2, for logistic models (2.4) (with $M=2$), the parameters $c,\gamma$, and threreby the transition weights $\alpha_{1,t},...,\alpha_{M,t}$$t=1,...,T$, are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$. With nonrandom exogenous transition weights, the identification of $\alpha_{1,t},...,\alpha_{M,t}$$t=1,...,T$ is given. In either case, the transition weights can be treated as uniquely identified in the rest of the proof, where, for readability, we do not repeatedly remind that the identification holds only almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$.
Denote by $\boldsymbol{\tau}_{m}=(\tau_{m1},...,\tau_{md})$ the permutation $\tau_{m1},...,\tau_{md}$ and $\boldsymbol{\iota}_{m}=(\iota_{m1},...,\iota_{md})$ the sign changes $\iota_{m1},...,\iota_{md}$ of the columns of $B_{m}$, respectively. Moreover, denote by $B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})$ the invertible impact matrix $B_{m}$ corresponding to the permutation $\boldsymbol{\tau}_{m}$ and sign changes $\boldsymbol{\iota}_{m}$ of its columns. Then, consider the permutations $\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{2},...,\boldsymbol{\tau}_{M}$ and $\boldsymbol{\tau}_{1},\tilde{\boldsymbol{\tau}}_{2},...,\tilde{\boldsymbol{\tau}}_{M}$ as well the sign changes $\boldsymbol{\iota}_{1},\boldsymbol{\iota}_{2},...,\boldsymbol{\iota}_{M}$ and $\boldsymbol{\iota}_{1},\tilde{\boldsymbol{\iota}}_{2},...,\tilde{\boldsymbol{\iota}}_{M}$. We start by showing that $B_{y,t}(\boldsymbol{\tau},\boldsymbol{\iota})\equiv\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})$ cannot be obtained by reordering or changing the signs of the columns of $B_{y,t}(\tilde{\boldsymbol{\tau}},\tilde{\boldsymbol{\iota}})\equiv\alpha_{m,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m})$ almost everywhere in $[B_{1},...,B_{M}]\in\mathbb{R}^{d\times Md}$, whenever $\boldsymbol{\tau}_{m}\neq\tilde{\boldsymbol{\tau}}_{m}$ or $\boldsymbol{\iota}_{m}\neq\tilde{\boldsymbol{\iota}}_{m}$ for any $m=2,...,M$.
@@
By the above discussion, for any single transformation $D_{t}P_{t}$ of $B_{y,t}$, the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (B.22) for any time period when all the entries of $(\alpha_{1,t},...,\alpha_{M,t})$ are nonzero has a Lebesgue measure zero in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$. Since there exists a finite number of such transformations and a countable union of measure zero sets has a measure zero, it follows that the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (B.23) for all $t$ has a Lebesgue measure zero in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$.
@@
Under Condition (C), the vector $(\alpha_{1,t},...,\alpha_{M,t})$ and all its subvectors $(\alpha_{n_{1},t},...,\alpha_{n_{M^{\circ}},t})$, $n_{l}\in\mathbb{M^{\circ}}$, $\mathbb{M^{\circ}}\subset\{1,...,M\}$, where $M^{\circ}\geq 2$ denotes the number of elements in $\mathbb{M^{\circ}}$, take at least $M+1$ linearly independent values (plus one accounting for transition weights that equal to zero). Therefore, Equation (B.25) defined together for all $t$ defines a system of linear equations involving $Md^{2}$ variables and at least $Md^{2}$ linearly independent equations. Moreover, as $(\alpha_{1,t},...,\alpha_{M,t})$ and all its subvectors take at least $M+1$ linearly independent values, each variable $\tilde{b}_{ij}^{(m)}$ appears in at least $M$ linearly independent the equations. Thus, there exists at most one solution to this system of linear equations. Since one solution is $\tilde{B}_{m}=B_{m}$ for all $m=1,...,M$, this is the unique solution, concluding that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywehere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, given that the transition weights $\alpha_{m,t}$, $m=1,...,M$, $t=1,...,T$, are uniquely identified. As already stated, the transition weights weights are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$ for logistic models and almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ for models with nonrandom exogenous weight, concluding the proof.∎